# It’s complicated: The relationship of algorithmic fairness and the non‑discrimination provisions for high‑risk systems in the EU AI Act

### v5

@@ 1 Introduction
How can we ensure fair Artificial Intelligence (AI) systems? When regulating the digital world, particularly AI systems, this question becomes fundamental. While the question is also crucial in a non-digital world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners in regulating the digital age. To improve the EU market and to promote the uptake of human-centric and trustworthy AI, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^1] Regulation (EU) 2024/1689., see Article 1 AIA. While the AIA is partly a product safety regulation (kaminski2025american; almada2023eu) — specifically, see Chapter III AIA — it also incorporates aspects of “ethical principles” (nolte2024robustness) within and outside the EU (see recital 27 AIA).
Fair algorithmic processing also matters from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating the discriminatory behaviour of Artificial Intelligence models (pessach2023algorithmic). The importance becomes evident when looking at several discriminatory behaviours by algorithms in the past, ranging from hiring (dastin2022amazon) to social welfare systems (hadwick2021lessons). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications, have posed significant new challenges to the algorithmic fairness domain (chu2024fairness; kotek2023gender).
@@ 1.1 Previous literature
Discrimination through algorithms is not a recent phenomenon and was analysed prior to the AI Act. As early as the 1980s, algorithmic discrimination was observed in admissions settings (connors1981national; williams1981analysis), and gender inequalities in educational software were discussed (huff1987sex). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (friedman1996bias). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. (barocas2016big) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied (wachter2021fairness; wachter2020bias; hacker2018teaching; weerts2023algorithmic; xenidis2020eu; lewis2025mapping). For example, the seminal study by wachter2021fairness presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analysed to “unlock the algorithmic black box” (hacker2018teaching). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
Recently, several papers have addressed the AI Act. Short studies such as (deck2024implications; ruohonen2024algorithmic) have analysed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (wachter2024limitations; hacker2024generative; novelli2024generative). correa2025better investigated fairness principles in the AI Act in the context of human oversight mechanisms. The paper of raman2025disclosure discussed the fairness provisions of the AI Act for General Purpose AI (GPAI) models. The AI Act and its relation to gender equality and non-discrimination law have also been discussed (lutz2024ai). Among these, the work by bosoer2023non is most similar to ours in terms of its focus on the AI Act; however, it investigates non-discrimination provisions in the draft of the AI Act without a strong focus on the interaction with computer science. The most similar work in terms of the interaction between computer science and legal research is the paper by weerts2023algorithmic. Unlike our work, theirs did not focus on the AI Act.
@@ 2 A primer on EU non-discrimination legislation and algorithmic fairness
However, the complexity of interdisciplinary work goes beyond technical expertise. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behaviour or outcomes of algorithms. Although fairness has been discussed as a principle in economic law contexts (scheuerer2023fairness), it is not a specific legal term. The closest legal term is arguably “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between these key concepts from law and computer science is demanding.
@@ 2.1 EU non-discrimination law: A gentle introduction
Fundamental rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is today one of the primary fundamental rights laws within the EU [^2] The CFR should not be confused with the European Convention on Human Rights (ECHR), which, though separate from the EU, is an agreement to which the EU is expected to accede. The relationship between the ECHR and EU law is complex (see brittain2015relationship).. Under Article 6(1) of the Treaty on European Union, the Charter forms part of the primary law of the EU.
@@
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). Interpretation and application of the Charter require balancing rights as described in Article 52(1) CFR. First, legislators must balance different CFR rights proportionately when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this proportionality requirement. A logic of proportionality guides the assessment of fundamental rights (almada2023eu). Fundamental rights are neither absolute, hierarchical, nor quantifiable, and must be applied on a case-by-case basis (sousa2024artificial). The provisions of the CFR apply to public parties (vertical applicability, fornasier2015impact). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (fornasier2015impact; cherednychenko2007fundamental; frantziou2015horizontal; prechal2020horizontal) and in the AI Act context specifically (lewis2025mapping). In the context of the GDPR, scholars and the European Court of Justice (ECJ) noted [^3] See case Case C-555/07, Kücükdeveci v Swedex GmbH & Co. KG, ECLI:EU:C:2010:21, para. 27. that the CFR can have a horizontal effect “when EU secondary law gives expression to a general principle of EU law, such as the principles of privacy and protection of personal data and non-discrimination.” (ufert2020ai). Therefore, some horizontal applicability of the CFR is assumed when applying the AI Act. When discussing fundamental rights in the AI Act, it is also essential to consider the broader political and institutional landscape (palmiotto2025ai).
@@
Direct and indirect discrimination in EU Law. While the CFR is part of EU primary law, it also influences more recent secondary EU law (lewis2025mapping). EU law has a long history of non-discrimination regulations [^4] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination_en.. For example, the Directive 2000/43/EC (Race Equality Directive) and Directive 2000/78/EC (Employment Equality Directive) include specific non-discrimination norms.
@@
In order to legally assess discrimination under specific secondary EU law, two types of discrimination are distinguished: direct and indirect discrimination (zuiderveen2024non; wachter2020bias). However, importantly, in EU law, direct discrimination does not require any intent to discriminate (weerts2023algorithmic; xenidis2020eu; adams2023directly).
@@
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC) [^5] In the cited directive, this targets discrimination on grounds of racial or ethnic origin.. This means that discrimination occurs when individuals are treated less favourably based on a protected attribute listed in the Article.
@@
Indirect discrimination is more difficult to address. In indirect discrimination cases, seemingly neutral attributes are used but they rely on a protected attribute. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (zuiderveen2024non).
@@
Also, a third category of discrimination needs to be taken into account [^6] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (xenidis2023computers). This becomes especially important since (modern) AI models do not use single variables as input but instead use many different aspects as input. This can lead to effects where discrimination may occur only at the intersection of gender and age (weerts2023algorithmic). Whether, and to what extent, the ECJ currently recognises intersectional discrimination as a distinct form of discrimination remains an open question(weerts2023algorithmic; xenidis2023computers; atrey2018illuminating).
@@ 2.2 Algorithmic fairness: A gentle introduction
Algorithmic fairness has emerged as an established computer science and AI-related field in recent years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (mitchell2021algorithmic). This treatment must be considered within appropriate social, theoretical, and legal contexts (pessach2023algorithmic) and includes evaluation and auditing methods to test for unfairness.
@@
Origins of unfairness in algorithms. Unfair algorithmic behaviour can have different sources. In order to discuss the origins, another — not well-defined — term from computer science is often used: bias. There exist many different bias definitions. Since the EU AI Act does not define the word bias itself, we use a definition from the European Commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (noauthor_commission_2021). Biases are primarily discussed within the data used for AI models and can be sorted into different categories. Most researchers emphasise the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (zou2018ai). (hacker2018teaching) categorises the biases in data into two groups: biased training data versus unequal ground truth. (barocas2016big) uses five categories to map biases, while (mehrabi2021survey) uses 21 different categories for biases. It is important to note that the data are only one — albeit important — source of biases. For example, design decisions in the algorithms themselves or structural power asymmetries also lead to algorithmic unfairness (mehrabi2021survey; gebru2021datasheets; sousa2024artificial).
@@
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (corbett2023measure; castelnovo2022clarification; verma2018fairness). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (dwork2012fairness) or groups of people (barocas2023fairness; binns2020apparent). They provide a quantitative measure intended to indicate whether an algorithm demonstrates unjustified unequal treatment. Fairness metrics also include moral norms (deck2024implications; hellman2020measuring).
@@
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on, and provide a rationale for the optimisation criteria chosen. First, it has been shown that an apparently objective number needs to be contextualized to the application under consideration (wachter2021fairness). Furthermore, some metrics are, from a mathematical viewpoint, incompatible with each other [^7] They are also incompatible from a moral point of view (heidari2019moral)., thus these fairness metrics cannot be satisfied at the same time (chouldechova2017fair; kleinberg2016inherent). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favourite metric to create the (technical and mathematical) impression that the algorithm is fair (black2024d; meding2024fairness). Fairness metrics for mitigation can play a role at different stages of the development, testing, and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are distinguished (barocas2023fairness; binns2018fairness). In the pre-processing phase, the input data itself is altered to ensure fair processing (caton2024fairness; kamiran2012data). For the in-training phase, the optimisation process during the training of an ML model is adjusted to include fairness as an optimisation goal. Finally, in the post-processing phase, the output of a pre-trained model is adapted. The first and last approach make it possible to perform fairness analyses even if one only has black box access to the model (caton2024fairness).
@@
The era of LLMs raises new questions. LLMs have the advantage — or disadvantage, depending on the viewpoint — that their exact use case is most of the time not predefined, and they are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, this diversity also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (doan2024fairness; chu2024fairness). Some researchers have applied classical fairness metrics in classification settings to LLMs (chhikara2024few). Additionally, it has been shown that LLM-specific issues arise due to the use of massive data and processing (kotek2023gender; ferrara2023should; navigli2023biases; huang2024trustllm) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (anthis2024impossibility).

### v6

@@ 1 Introduction
How can we ensure fair Artificial Intelligence (AI) systems? When regulating the digital world, particularly AI systems, this question becomes fundamental. While the question is also crucial in a non-digital world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners in regulating the digital age. To improve the EU market and to promote the uptake of human-centric and trustworthy AI, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^1] Regulation (EU) 2024/1689., see Article 1 AIA. While the AIA is partly a product safety regulation (Kaminski and Selbst, 2025; Almada and Petit, 2023) — specifically, see Chapter III AIA — it also incorporates aspects of “ethical principles” (Nolte et al., 2024) within and outside the EU (see recital 27 AIA).
Fair algorithmic processing also matters from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating the discriminatory behaviour of Artificial Intelligence models (Pessach and Shmueli, 2023). The importance becomes evident when looking at several discriminatory behaviours by algorithms in the past, ranging from hiring (Dastin, 2022) to social welfare systems (Hadwick and Lan, 2021). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications, have posed significant new challenges to the algorithmic fairness domain (Chu et al., 2024; Kotek et al., 2023).
@@ 1.1 Previous literature
Discrimination through algorithms is not a recent phenomenon and was analysed prior to the AI Act. As early as the 1980s, algorithmic discrimination was observed in admissions settings (Connors et al., 1981; Williams et al., 1981), and gender inequalities in educational software were discussed (Huff and Cooper, 1987). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (Friedman and Nissenbaum, 1996). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. (Barocas and Selbst, 2016) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied (Wachter et al., 2021, 2020; Hacker, 2018; Weerts et al., 2023; Xenidis and Senden, 2020; Lewis et al., 2025). For example, the seminal study by Wachter et al. (2021) presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analysed to “unlock the algorithmic black box” (Hacker, 2018). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
Recently, several papers have addressed the AI Act. Short studies such as (Deck et al., 2024a; Ruohonen, 2024) have analysed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (Wachter, 2024; Hacker et al., 2024b; Novelli et al., 2024). Corrêa et al. (2025) investigated fairness principles in the AI Act in the context of human oversight mechanisms. The paper of Raman et al. (2025) discussed the fairness provisions of the AI Act for General Purpose AI (GPAI) models. The AI Act and its relation to gender equality and non-discrimination law have also been discussed (Lütz, 2024). Among these, the work by Bosoer et al. (2023) is most similar to ours in terms of its focus on the AI Act; however, it investigates non-discrimination provisions in the draft of the AI Act without a strong focus on the interaction with computer science. The most similar work in terms of the interaction between computer science and legal research is the paper by Weerts et al. (2023). Unlike our work, theirs did not focus on the AI Act.
@@ 2 A primer on EU non-discrimination legislation and algorithmic fairness
However, the complexity of interdisciplinary work goes beyond technical expertise. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behaviour or outcomes of algorithms. Although fairness has been discussed as a principle in economic law contexts (Scheuerer, 2023), it is not a specific legal term. The closest legal term is arguably “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between these key concepts from law and computer science is demanding.
@@ 2.1 EU non-discrimination law: A gentle introduction
Fundamental rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is today one of the primary fundamental rights laws within the EU [^2] The CFR should not be confused with the European Convention on Human Rights (ECHR), which, though separate from the EU, is an agreement to which the EU is expected to accede. The relationship between the ECHR and EU law is complex (see Brittain (2015)).. Under Article 6(1) of the Treaty on European Union, the Charter forms part of the primary law of the EU.
@@
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). Interpretation and application of the Charter require balancing rights as described in Article 52(1) CFR. First, legislators must balance different CFR rights proportionately when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this proportionality requirement. A logic of proportionality guides the assessment of fundamental rights (Almada and Petit, 2023). Fundamental rights are neither absolute, hierarchical, nor quantifiable, and must be applied on a case-by-case basis (Sousa e Silva, 2024). The provisions of the CFR apply to public parties (vertical applicability, Fornasier (2015)). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (Fornasier, 2015; Cherednychenko, 2007; Frantziou, 2015; Prechal, 2020) and in the AI Act context specifically (Lewis et al., 2025). In the context of the GDPR, scholars and the European Court of Justice (ECJ) noted [^3] See case C-555/07, Kücükdeveci, para 27. that the CFR can have a horizontal effect “when EU secondary law gives expression to a general principle of EU law, such as the principles of privacy and protection of personal data and non-discrimination.” (Ufert, 2020). Therefore, some horizontal applicability of the CFR is assumed when applying the AI Act. When discussing fundamental rights in the AI Act, it is also essential to consider the broader political and institutional landscape (Palmiotto, 2025).
@@
Direct and indirect discrimination in EU Law. While the CFR is part of EU primary law, it also influences more recent secondary EU law (Lewis et al., 2025). EU law has a long history of non-discrimination regulations [^4] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination_en.. For example, the Directive 2000/43/EC (Race Equality Directive) and Directive 2000/78/EC (Employment Equality Directive) include specific non-discrimination norms.
@@
In order to legally assess discrimination under specific secondary EU law, two types of discrimination are distinguished: direct and indirect discrimination (Borgesius et al., 2024; Wachter et al., 2020). However, importantly, in EU law, direct discrimination does not require any intent to discriminate (Weerts et al., 2023; Xenidis and Senden, 2020; Adams-Prassl et al., 2023).
@@
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC) [^5] In the cited directive, this targets discrimination on grounds of racial or ethnic origin. This means that discrimination occurs when individuals are treated less favourably based on a protected attribute listed in the Article.
@@
Indirect discrimination is more difficult to address. In indirect discrimination cases, seemingly neutral attributes are used but they rely on a protected attribute. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (Borgesius et al., 2024).
@@
Also, a third category of discrimination needs to be taken into account [^6] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis, 2023). This becomes especially important since (modern) AI models do not use single variables as input but instead use many different aspects as input. This can lead to effects where discrimination may occur only at the intersection of gender and age (Weerts et al., 2023). Whether, and to what extent, the ECJ currently recognises intersectional discrimination as a distinct form of discrimination remains an open question(Weerts et al., 2023; Xenidis, 2023; Atrey, 2018).
@@ 2.2 Algorithmic fairness: A gentle introduction
Algorithmic fairness has emerged as an established computer science and AI-related field in recent years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (Mitchell et al., 2021). This treatment must be considered within appropriate social, theoretical, and legal contexts (Pessach and Shmueli, 2023) and includes evaluation and auditing methods to test for unfairness.
@@
Origins of unfairness in algorithms. Unfair algorithmic behaviour can have different sources. In order to discuss the origins, another — not well-defined — term from computer science is often used: bias. There exist many different bias definitions. Since the EU AI Act does not define the word bias itself, we use a definition from the European Commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (Commission, 2021). Biases are primarily discussed within the data used for AI models and can be sorted into different categories. Most researchers emphasise the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (Zou and Schiebinger, 2018). (Hacker, 2018) categorises the biases in data into two groups: biased training data versus unequal ground truth. (Barocas and Selbst, 2016) uses five categories to map biases, while (Mehrabi et al., 2021) uses 21 different categories for biases. It is important to note that the data are only one — albeit important — source of biases. For example, design decisions in the algorithms themselves or structural power asymmetries also lead to algorithmic unfairness (Mehrabi et al., 2021; Gebru et al., 2021; Sousa e Silva, 2024).
@@
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (Corbett-Davies et al., 2023; Castelnovo et al., 2022; Verma and Rubin, 2018). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (Dwork et al., 2012) or groups of people (Barocas et al., 2023; Binns, 2020). They provide a quantitative measure intended to indicate whether an algorithm demonstrates unjustified unequal treatment. Fairness metrics also include moral norms (Deck et al., 2024a; Hellman, 2020).
@@
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on, and provide a rationale for the optimisation criteria chosen. First, it has been shown that an apparently objective number needs to be contextualized to the application under consideration (Wachter et al., 2021). Furthermore, some metrics are, from a mathematical viewpoint, incompatible with each other [^7] They are also incompatible from a moral point of view (Heidari et al., 2019)., thus these fairness metrics cannot be satisfied at the same time (Chouldechova, 2017; Kleinberg et al., 2016). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favourite metric to create the (technical and mathematical) impression that the algorithm is fair (Black et al., 2024; Meding and Hagendorff, 2024). Fairness metrics for mitigation can play a role at different stages of the development, testing, and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are distinguished (Barocas et al., 2023; Binns, 2018). In the pre-processing phase, the input data itself is altered to ensure fair processing (Caton and Haas, 2024; Kamiran and Calders, 2012). For the in-training phase, the optimisation process during the training of an ML model is adjusted to include fairness as an optimisation goal. Finally, in the post-processing phase, the output of a pre-trained model is adapted. The first and last approach make it possible to perform fairness analyses even if one only has black box access to the model (Caton and Haas, 2024).
@@
The era of LLMs raises new questions. LLMs have the advantage — or disadvantage, depending on the viewpoint — that their exact use case is most of the time not predefined, and they are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, this diversity also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (Doan et al., 2024; Chu et al., 2024). Some researchers have applied classical fairness metrics in classification settings to LLMs (Chhikara et al., 2024). Additionally, it has been shown that LLM-specific issues arise due to the use of massive data and processing (Kotek et al., 2023; Ferrara, 2023; Navigli et al., 2023; Huang et al., 2024) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (Anthis et al., 2024).